OpenStudy (anonymous):

if x=3 and y=-2 then x^2 + 2y =?

OpenStudy (anonymous):

Plug the values in my friend!!

OpenStudy (anonymous):

3 ^ 2 + 2 (-2)

OpenStudy (diyadiya):

3^2+2(-2)=9-4=5